2. An ultracentrifuge accelerates from rest to 100,000 rpm in 2.00 min. (a) What is its angular acceleration in rad/s2? (b) What is the tangential acceleration of a point 9.50 cm from the axis of rotation?
1
Expert's answer
2021-10-13T15:42:22-0400
"\alpha=\frac{\pi}{30}\frac{100000}{120}=87.3\frac{rad}{s^2}\\a_t=87.3(0.095)=8.29\frac{m}{s^2}\\"